OpenStudy (anonymous):

in how many ways can we divide 5 different coloured balls to three persons such that everyone must get atleast one ball..

OpenStudy (anonymous):

Refer http://en.wikipedia.org/wiki/Stars_and_bars_%28combinatorics%29 and get your answer

OpenStudy (anonymous):

but that one will apply only to identical balls..

OpenStudy (anonymous):

Similar problem asked by me http://openstudy.com/study#/updates/4f95edc7e4b000ae9ecbd6eb

OpenStudy (blockcolder):

First give each person a ball. This can be done in 5P3 ways. Then distribute the remaining two balls to the 3 persons. This can be done is \(\large \binom{3}{2}\cdot 2\) ways. Not entirely sure, though. If there are any errors in the reasoning, feel free to point them out.

OpenStudy (anonymous):

Oops. Read stirling number of 2nd kind to get your solution

OpenStudy (anonymous):

Sorry @quarkine for telling you to refer Star and bar Combinatorics

OpenStudy (anonymous):

@shivam_bhalla think maybe i can couple the star one with 5! to get the ans...@blockcolder :lemme check ...

OpenStudy (anonymous):

@quarkine: S(5,3)

OpenStudy (anonymous):

@blockcolder :i think you havent considered the possibility that the remaining all two balls can be given to just one of the three..

OpenStudy (anonymous):

You will also need 3!. now you can join them somehow to get the answer ;)

OpenStudy (anonymous):

it looks like the balls can not be less than 1 in this case there are 5 ways in which one person can get balls. only 1 ball . that means 5 ways for a ball, then (5+ 5p2+5p3+5p4+5factorial) This is for one person. multiply it by 3 haha i am getting 975 ways.

OpenStudy (phi):

Number of ways to assign 3 balls to person A, and 1 each to persons B and C \[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\left(\begin{matrix}2 \\ 1\end{matrix}\right)\left(\begin{matrix}1 \\ 1\end{matrix}\right)\] There are 3C1 ways to choose the person to get the 3 balls So there are 10*2*1* 3= 60 ways for one person to have 3 balls and the other 2 have 1 ball. There are \[\left(\begin{matrix}5 \\ 2\end{matrix}\right)\left(\begin{matrix}3 \\ 2\end{matrix}\right)\left(\begin{matrix}1 \\ 1\end{matrix}\right)\] ways for person A to have 2 balls, person B to have 2 balls and C to have 1. There are 3C1 ways to choose the person to get 1 ball. So 10*3*1 *3 = 90 ways for 2,2,1 Total number of ways= 150

OpenStudy (anonymous):

@FoolForMath :no i dont hav the answer...

OpenStudy (anonymous):

i think it may be 5P3, as @blockcolder said, times 3X3(which would take care of the possible ways to distribute the remaining 2) so it should be 540..

OpenStudy (anonymous):

anyone got the answer to this??

OpenStudy (anonymous):

Have a look at @phi 's answer